# Wide Neural Networks as a Baseline for the Computational No-Coincidence Conjecture

### v1

@@ 1 Introduction
The zero-mean criterion is satisfied for tanh but not ReLU or GeLU functions, agreeing with prior research that suggests in different ways that tanh networks are biased towards complex functions \@BBOPcitep\@BAP\@BBN(Poole et al., 2016)\@BBCP, and ReLU networks are biased towards simple functions \@BBOPcitep\@BAP\@BBN(Hanin & Rolnick, 2019; Palma et al., 2019; Teney et al., 2024)\@BBCP. But the criterion can also be satisfied by any activation function with an additive shift, such as $\sigma(z)=\text{ReLU}(z)-1/\sqrt{2\pi}$.
@@ 2 Problem Setup
Here, $\sigma:\mathbb{R}\to\mathbb{R}$ is our activation function, $n$ is our network width, and at every layer $\ell$, our networks are parameterized by the weight matrix $\{W^{(\ell)}_{ij}\}_{i,j=1,\ldots,n}$ and bias vector $\{b^{(\ell)}_{i}\}_{i=1,\ldots,n}$. Instead of a dedicated output layer, we consider the preactivations at an arbitrary layer as the final outputs of our neural networks. Our weights and biases are drawn from a Gaussian distribution with a mean of zero and variance set by constants $C_{b}^{(\ell)}$ and $C_{W}^{(\ell)}$:
@@ 4.1 Hermite Decompositions
To understand (16) and the behavior of $\mathcal{C}$, we can decompose our activation function $\sigma$ into Hermite polynomials $\textrm{He}_{n}$ which are orthogonal under Gaussian weighting \@BBOPcitep\@BAP\@BBN(Bateman & Erdélyi, 1953)\@BBCP.
@@ 5 Conclusion
For a randomly initialized neural network $C:\mathbb{R}^{n}\to\mathbb{R}^{n}$ with hidden dimension $n$, depth $\ell=\Theta(\log(n))$, Gaussian random weights, and tanh nonlinearities, let $P(C)$ be the property that for none of the inputs in the dataset $x\in\{-1,1\}^{n}$, the output $C(x)$ is all negative. There exists a polynomial-time verification algorithm $V$ that receives as input the neural network $C$ and an advice string $\pi$ such that for all $C$ where $P(C)$ is true, there exists $\pi$ with length polynomial in the size of $C$ such that $V(C,\pi)=1$, and for 99% of random neural networks $C$, no such $\pi$ exists.

### v2

@@ 1 Introduction
The zero-mean criterion is satisfied for tanh but not ReLU or GeLU functions, agreeing with prior research that suggests in different ways that tanh networks are biased towards complex functions \@BBOPcitep\@BAP\@BBN(Poole et al., 2016)\@BBCP, and ReLU networks are biased towards simple functions \@BBOPcitep\@BAP\@BBN(Hanin and Rolnick, 2019; Palma et al., 2019; Teney et al., 2024)\@BBCP. But the criterion can also be satisfied by any activation function with an additive shift, such as $\sigma(z)=\text{ReLU}(z)-1/\sqrt{2\pi}$.
@@ 2 Problem Setup
Here, $\sigma:\mathbb{R}\to\mathbb{R}$ is our activation function, $n$ is our network width, and at every layer $\ell$, our networks are parameterized by the weight matrix $\{W^{(\ell)}_{ij}\}_{i,j=1,\ldots,n}$ and bias vector $\{b^{(\ell)}_{i}\}_{i=1,\ldots,n}$. Instead of a dedicated output layer, we consider the preactivations at an arbitrary layer as the final outputs of our neural networks, and for outputs smaller than the network width, we can freely ignore the extra neurons. Our weights and biases are drawn from a Gaussian distribution with a mean of zero and variance set by constants $C_{b}^{(\ell)}$ and $C_{W}^{(\ell)}$:
@@ 4.1 Hermite Decompositions
To understand (16) and the behavior of $\mathcal{C}$, we can decompose our activation function $\sigma$ into Hermite polynomials $\textrm{He}_{n}$ which are orthogonal under Gaussian weighting \@BBOPcitep\@BAP\@BBN(Bateman and Erdélyi, 1953)\@BBCP.
@@ 5 Conclusion
For a randomly initialized neural network $C:\mathbb{R}^{2n}\to\mathbb{R}^{n}$ with hidden dimension $2n$, depth $\ell=\Theta(\log(n))$, Gaussian random weights, and tanh nonlinearities, let $P(C)$ be the property that for none of the inputs in the dataset $x\in\{-1,1\}^{2n}$, the output $C(x)$ is all negative. There exists a polynomial-time verification algorithm $V$ that receives as input the neural network $C$ and an advice string $\pi$ such that for all $C$ where $P(C)$ is true, there exists $\pi$ with length polynomial in the size of $C$ such that $V(C,\pi)=1$, and for 99% of random neural networks $C$, no such $\pi$ exists.